What is the probability of picking a 7 out of a deck of cards?
Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

4/52=.0769=8% chance

Answer by stanbon(75887)

What is the probability of picking a 7 out of a deck of cards?
--------------------
# of ways to succeed: 4
# of random outcomes:52
P(7) = 4/52 = 1/13
=====================
Cheers,
Stan H.